If a body is released from certain height above the ground in the last second of its fall it covers half of the total distance traveled by it then total time of its fall is

Grade:

1 Answers

Arun
5 years ago

As we know that,

S = ut+½at²

u = 0 ( freefall )

Let total time be ‘t’

Total distance be ‘S’

So,

S/2 = ½(9.8)(T-1)² -[1]

S = ½(9.8)(T)² -[2]

From [1] & [2]

2{½(9.8)(T-1)²} = ½(9.8)(T)²

2(T-1)² = (T)²

√2(T-1) = T

T= √2/(√2–1)

T = 3.4142 second.